# Quantifying non-Hermiticity using single- and many-particle quantum properties

### v1

@@ II.3.1 Hamiltonian non-Hermiticity
$\mathcal{D}=\frac{||H_{\mathrm{nh}}-H_{\mathrm{nh}}^{\dagger}||}{||H_{\mathrm{nh}}||}\,,$ (13)
where ||.||||.||| | . | | is the operator norm. This measure quantifies the distance between a model Hamiltonian $H_{\mathrm{nh}}$ and its Hermitian conjugate. It naturally vanishes for Hermitian systems. We find it most useful to normalize $\mathcal{D}$ to the norm of $H_{\mathrm{nh}}$, in order to avoid attributing significant non-Hermiticity to systems that are actually dominated by a large Hermitian part (see also Appendix C for a discussion about the unnormalized version). While it captures properties that the entire model system can in principle assume, quantifying non-Hermiticity via $\mathcal{D}$ requires access to the full Hamiltonian matrix.
II.3.2 Non-Hermiticity score of observables
@@
This score quantifies the non-Hermiticity at the level of any computable physical quantity related to the model. It measures the difference of a state-dependent function $\mathcal{F}[.]$ when evaluated for the state evolved under $H_{\mathrm{nh}}$ or under $H_{\mathrm{nh}}^{\dagger}$, which as discussed above lead to ensembles of right or left eigenvectors ($\rho_{RR}(t)$ or $\rho_{LL}(t)$), respectively. As illustrated in the next section, one can choose for $\mathcal{F}[.]$ single-body observables such as local magnetization or occupations as well as many-body observables such as von-Neumann entropy or purity. Following the discussions in Sec. II.1, the corresponding expectation values are well-defined. The non-Hermiticity score thus expresses how much the non-Hermiticity of the model Hamiltonian is concretely translated into physical observables in a given time evolution.
Following also the discussion in Sec. II.2, the density matrices $\rho_{RR}$ and $\rho_{LL}$ used to evaluate the score can represent the snapshot of a system at a given evolution time (see also Fig. 3 below) or they can correspond to individual eigenstates (Figs. 2). For the latter case, sometimes the non-Hermiticity score can depend significantly on the choice of eigenvectors. To have a state-independent analysis that captures the behavior across the spectrum, we can reformulate the non-Hermiticity score of Eq. (14) as a well-defined distance measure, given by
$||\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}||_{p}=\Big{(}\sum_{i=1}\big{(}{{SC}^{\mathcal{F}}_{\mathrm{nh}}}[i]\big{)}^{p}\Big{)}^{\frac{1}{p}}.$ (15)
This quantity is nothing but the $p$-norm of the vector $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}$ with elements $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}=\{{SC}^{\mathcal{F}}_{\mathrm{nh}}[1],{SC}^{\mathcal{F}}_{\mathrm{nh}}[2],\dots,{SC}^{\mathcal{F}}_{\mathrm{nh}}[D]\}$, where ${SC}^{\mathcal{F}}_{\mathrm{nh}}[k]$ denotes the non-Hermiticity score computed for the $k$-th eigenstate of the model, and $D$ corresponds to the Hilbert space dimension. In our case, we will illustrate the results obtained for the case $p\rightarrow\infty$ (see Fig. 4), for which the above definition reduces to the infinity norm, $||\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}||_{\infty}=\max_{i}{SC}^{\mathcal{F}}_{\mathrm{nh}}[i]$, which is the maximum value of the non-Hermiticity score computed over all the eigenstates of the model.
At the opposite limit, $p\rightarrow 0^{+}$, one obtains the 00-norm, which counts the number of eigenvectors with non-zero non-Hermiticity score and thus captures the qualitative behavior across the entire spectrum. In our analysis below, for convenience we use a numerical threshold $\mathcal{E}_{\mathrm{T}h}^{\mathcal{F}}$, counting the total number of elements of $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}$ with the condition ${SC}^{\mathcal{F}}_{\mathrm{nh}}[k]\geq\mathcal{E}_{\mathrm{T}h}^{\mathcal{F}}$. We denote the corresponding score by $\mathcal{G}^{\mathcal{F}}$. An example is plotted in Fig. 5.
@@ III.1 Non-Hermitian model construed using imperfect Bell state
In order to introduce the first example of the non-Hermitian Hamiltonian that we have considered in our work, we present a set of “imperfect” Bell states characterized by the parameter $\alpha$ ($10<\alpha<$ 1) and having the following mathematical form when expressed in the computational basis, given by
@@
The non-Hermitian Hamiltonian ($H^{\mathrm{Bell}}_{\mathrm{nh}}$) is constructed by choosing the imperfect Bell states as the set of right eigenvectors. Subsequently, the set of left eigenvectors can be obtained from the inverse of a matrix whose columns are the vectors $\left\{|R_{m}\rangle\right\}$. The unnormalized left eigenvectors are
@@
$H^{\mathrm{Bell}}_{\mathrm{nh}}=\sum_{m}\lambda_{m}|R_{m}\rangle\langle L_{m}|\,,$ (19)
@@ III.1.1 Non-Hermiticity at the level of Hamiltonian:
We first examine how the degree of non-Hermiticity $\mathcal{D}$ behaves for the Hamiltonian described in Eq. (19). For a fixed value of the parameter $\alpha$, we compute $\mathcal{D}$ for 1000 random realizations of $H_{\mathrm{nh}}$ by choosing its eigenvalues $\{\lambda_{m}\}$ from a Gaussian distribution with zero mean and unit variance. In Fig. 1(a), we plot the maximum as well as minimum among all the random realizations, denoted by $\mathcal{D}_{\max}$ and $\mathcal{D}_{\min}$, respectively, against $\bar{\alpha}=1-\alpha$. At $\alpha=1$ ($\bar{\alpha}=0$), the right and left eigenvectors recover perfect Bell states and the Hamiltonian becomes Hermitian, reflected in vanishing $\mathcal{D}$. Though fluctuating, $\mathcal{D}_{\min}$ tends to increase with $\bar{\alpha}$. The growth of $\mathcal{D}_{\max}$ is smoother and at a faster rate, reaching the maximum non-Hermiticity of $\mathcal{D}=1$ at $\bar{\alpha}=1$.
@@ III.1.2 Non-Hermiticity score for single-site property: magnetization
One of the main goals of our work is to examine to what extent the non-Hermiticity of the model Hamiltonian gets translated to its physical properties. Towards that aim, we now analyze the non-Hermiticity score for the magnetization in individual eigenstates. Results for quantum entanglement in individual eigenstates are presented in the following section. Analytical forms of all the quantities discussed in these two sections are provided in Appendix A. Two subsequent sections contain results for purity and entanglement under time evolution.
In Fig. 2(a), we plot the non-Hermiticity score for the magnetization of the first qubit (labeled $A$) along the $z$-axis, $\mathcal{F}[\rho]=m_{z}[\rho]=\text{Tr}(\hat{\sigma}^{z}_{A}\rho)$, $SC^{m_{z}}_{k}=|m_{z}[\rho_{RR}^{k}]-m_{z}[\rho_{LL}^{k}]|$ Here, $\hat{\sigma}^{z}$ is the $2\times 2$ Pauli matrix along the $z$-direction and $\rho_{RR}^{k}=\frac{|R_{k}\rangle\langle R_{k}|}{\langle R_{k}|R_{k}\rangle}$, $\rho_{LL}^{k}=\frac{|L_{k}\rangle\langle L_{k}|}{\langle L_{k}|L_{k}\rangle}$, where $|R_{k}\rangle$ and $|L_{k}\rangle$ are individual right and left eigenvectors of $H^{\mathrm{Bell}}_{\mathrm{nh}}$, as given in Eqs. (16) and (17). As we are considering here properties of single eigenvectors, the eigenvalues $\lambda_{k}$ do not directly enter the non-Hermiticity score, in contrast to $\mathcal{D}$ discussed above, and no sampling over $\lambda_{k}$ is necessary. To give an operational interpretation of the above observables, we may want to assume the eigenvalues to fulfill case 2 in Sec. II.2, such that $\rho_{RR}^{k}$ and $\rho_{LL}^{k}$ are prepared when evolving an initial state to long times under $H^{\mathrm{Bell}}_{\mathrm{nh}}$ and $(H^{\mathrm{Bell}}_{\mathrm{nh}})^{\dagger}$, respectively.
@@ III.1.3 Non-Hermiticity score for the von Neumann entropy (static case)
In this section, we analyze entanglement properties of the same set of right and left eigenvectors as considered above, in particular of the von Neumann entropy (VNE), defined as $\mathcal{S}[\rho]=-\text{Tr}_{A}(\text{Tr}_{B}(\rho)\log_{2}\text{Tr}_{B}(\rho))$. Figure 2(b) presents the associated non-Hermiticity score, $SC^{\mathcal{\mathcal{S}}}_{k}=|\mathcal{S}[\rho_{RR}^{k}]-\mathcal{S}[\rho_{LL}^{k}]|$. Unlike the case for $m_{z}$, at each value of $\bar{\alpha}$, $SC^{\mathcal{S}}_{k}$ becomes minimum for $k=1,2$, and the behavior does not comply with $\mathcal{D}$. In particular, $SC^{\mathcal{S}}_{k=1,2}$ vanishes both for $\bar{\alpha}\rightarrow 0$ and $\bar{\alpha}\rightarrow 1$, and attains a maximum around $\bar{\alpha}=0.449$. This is significantly different from the behavior of $\mathcal{D}_{\max}$ (or $\mathcal{D}_{\min}$) as well as the single-site property discussed above.
In contrast, the maximum value of $SC^{\mathcal{S}}_{k}$, attained for all values of $\bar{\alpha}$ at $k=3,4$, remains akin to $\mathcal{D}$. Nevertheless, quantitative differences remain. E.g., unlike $\mathcal{D}$, $SC^{\mathcal{S}}_{k=3,4}$ increases almost linearly in the region $0.4\lesssim\bar{\alpha}\leq 1.0$ and goes to zero at $\bar{\alpha}=0$. Moreover, $SC^{\mathcal{S}}_{k=3,4}$ does not reach its theoretical maximum of $1$ at $\bar{\alpha}=1$. Again, the VNE obtained for the biorthogonal ensemble $\rho_{RL}^{k}=|R_{k}\rangle\langle L_{k}|$ remains independent of $\bar{\alpha}$ (yielding the value $\mathcal{S}(|R_{k}\rangle\langle L_{k}|)=\frac{1}{2}$ $\forall k$) and is thus not a useful probe for the non-Hermiticity of the system.
@@ III.1.4 Non-Hermiticity score of global purity
In this section, we study the time-dependent non-Hermiticity score of the purity, $\mathcal{P}[\rho(t)]={\text{Tr}\Big{(}\rho(t)^{2}\Big{)}}/{\text{Tr}(\rho(t))^{2}}$, of a state $\rho(t)$ obtained from an initial state $\rho_{\mathrm{in}}$ when evolved under $H^{\mathrm{Bell}}_{\mathrm{nh}}$ or $(H^{\mathrm{Bell}}_{\mathrm{nh}})^{\dagger}$. If the initial state is chosen as maximally mixed, $\rho_{\mathrm{in}}=\frac{\mathbb{I}}{4}$, one gets $\mathcal{P}(\rho_{RR}(t))=\mathcal{P}(\rho_{LL}(t))$ (see Appendix B). Hence, $SC^{\mathcal{P}}(t)=0$ at all times and fails to witness any imprint of $\mathcal{D}$. However, for a non-maximally mixed initial state, $SC^{\mathcal{P}}(t)$ acquires a non-zero value. In Fig. 3, we plot the time-dependent purity score for systems initialized in a mixed Werner state, $\rho_{W}=\delta|\Psi^{-}\rangle\langle\Psi^{-}|+(1-\delta)\frac{\mathbb{I}}{4}$, with $|\Psi^{-}\rangle=\frac{|01\rangle-|10\rangle}{\sqrt{2}}$. We tune the initial purity to (a) $\delta=0.1$, (b) $\delta=0.5$, and (c) $\delta=0.9$, and choose four different values of $\alpha$, $\alpha=0.1$, $0.3$, $0.6$, $0.9$, which are depicted using lighter to darker shades of blue. In these plots, we fix the Hamiltonian eigenvalues to $\lambda_{1}=0.1$, $\lambda_{2}=0.2$, $\lambda_{3}=0.3$, and $\lambda_{4}=0.4$.
@@
Figure 3: Behavior of non-Hermiticity score with time, for global purity $\mathcal{P}$ and von Neumann entropy $\mathcal{S}$, computed for an initial state $\rho_{W}=\delta|\Psi^{-}\rangle\langle\Psi^{-}|+(1-\delta)\frac{\mathbb{I}}{4}$ subjected to evolution under $H^{\mathrm{Bell}}_{\mathrm{nh}}$ and $(H^{\mathrm{Bell}}_{\mathrm{nh}})^{\dagger}$. We probe the cases for $\alpha$=0.1, 0.3, 0.6, 0.9 as shown using curves with darker to lighter shades of blue, and in all cases fix the Hamiltonian eigenvalues to $\lambda_{1}=0.1,\ \lambda_{2}=0.2,\ \lambda_{3}=0.3,\ \lambda_{4}=0.4$. (a)-(c) Purity score, $SC^{\mathcal{P}}(t)$ for (a) $\delta=0.1$, (b) $\delta=0.5$, and (c) $\delta=0.9$. For large $\delta$, the amplitude of $SC^{\mathcal{P}}(t)$ decreases with increasing $\alpha$, while for small $\delta$ it attains a maximum at intermediate values of $\alpha$. (d)-(f) Non-Hermiticity score for the VNE, $SC^{\mathcal{S}}(t)$, for same choices of $\delta$.

### v2

@@ II.0.1 Hamiltonian non-Hermiticity
[…]
This measure quantifies the distance between a model Hamiltonian $\hat{H}_{\mathrm{nh}}$ and its Hermitian conjugate. It naturally vanishes for Hermitian systems. We find it most useful to normalize $\mathcal{D}$ to the norm of $\hat{H}_{\mathrm{nh}}$, in order to avoid attributing significant non-Hermiticity to systems that are actually dominated by a large Hermitian part (see also Appendix B for a discussion about the unnormalized version). While it captures properties that the entire model system can in principle assume, quantifying non-Hermiticity via $\mathcal{D}$ requires access to the full Hamiltonian matrix.
II.0.2 Non-Hermiticity score of observables
@@
This score quantifies the non-Hermiticity at the level of any computable physical quantity related to the model. It measures the difference of a state-dependent function $\mathcal{F}[.]$ when evaluated for the state evolved under $\hat{H}_{\mathrm{nh}}$ or under $\hat{H}_{\mathrm{nh}}^{\dagger}$, which as discussed above lead to ensembles of right or left eigenvectors ($\rho_{RR}(t)$ or $\rho_{LL}(t)$), respectively. As illustrated in the next section, one can choose for $\mathcal{F}[.]$ single-body observables such as local magnetization or occupations as well as many-body observables such as von-Neumann entropy or purity. Following the discussions in Sec. F, the corresponding expectation values are well-defined. The non-Hermiticity score thus expresses how much the non-Hermiticity of the model Hamiltonian is concretely translated into physical observables in a given time evolution.
Following also the discussion in Appendix C, the density matrices $\rho_{RR}$ and $\rho_{LL}$ used to evaluate the score can represent the snapshot of a system at a given evolution time (see also Fig. 3 below) or they can correspond to individual eigenstates (Figs. 2). For the latter case, sometimes the non-Hermiticity score can depend significantly on the choice of eigenvectors. To have a analysis that captures the behavior across the spectrum, we can reformulate the non-Hermiticity score of Eq. (2) as a well-defined distance measure, given by
[…]
This quantity is nothing but the $p$-norm of the vector $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}$ with elements $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}=\{{SC}^{\mathcal{F}}_{\mathrm{nh}}[1],{SC}^{\mathcal{F}}_{\mathrm{nh}}[2],\dots,{SC}^{\mathcal{F}}_{\mathrm{nh}}[D]\}$, where ${SC}^{\mathcal{F}}_{\mathrm{nh}}[k]$ denotes the non-Hermiticity score computed for the $k$-th eigenstate of the model, and $D$ corresponds to the Hilbert space dimension. In our case, we will illustrate the results obtained for the case $p\rightarrow\infty$ (see Fig. 4), for which the above definition reduces to the infinity norm, $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}||_{\infty}=\max_{i}{SC}^{\mathcal{F}}_{\mathrm{nh}}[i]$, which is the maximum value of the non-Hermiticity score computed over all the eigenstates of the model.
At the opposite limit, $p\rightarrow 0^{+}$, one obtains the 0-norm, which counts the number of eigenvectors with non-zero non-Hermiticity score and thus captures the qualitative behavior across the entire spectrum. In our analysis below, for convenience we use a numerical threshold $\mathcal{E}_{\mathrm{T}h}^{\mathcal{F}}$, counting the total number of elements of $\overline{{SC}^{\mathcal{F}}_{\mathrm{nh}}}$ with the condition ${SC}^{\mathcal{F}}_{\mathrm{nh}}[k]\geq\mathcal{E}_{\mathrm{T}h}^{\mathcal{F}}$. We denote the corresponding score by $\mathcal{G}^{\mathcal{F}}$. An example is plotted in Fig. 5.
@@ III.1 Non-Hermitian model construed using imperfect Bell state
In order to introduce the first example of the non-Hermitian Hamiltonian that we have considered in our work, we present a set of “imperfect” Bell states characterized by the parameter $\alpha$ ($0<\alpha\leq 1$) and having the following mathematical form when expressed in the computational basis, given by
@@
The non-Hermitian Hamiltonian ($\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}}$) is constructed by choosing the imperfect Bell states as the set of right eigenvectors. Subsequently, the set of left eigenvectors can be obtained from the inverse of a matrix whose columns are the vectors $\left\{|R_{m}\rangle\right\}$. The unnormalized left eigenvectors are
@@
$\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}}=\sum_{m}\lambda_{m}|R_{m}\rangle\langle L_{m}|\,$, (7)
@@ III.1.1 Non-Hermiticity at the level of Hamiltonian:
We first examine how the degree of non-Hermiticity $\mathcal{D}$ behaves for the Hamiltonian described in Eq. (7). For a fixed value of the parameter $\alpha$, we compute $\mathcal{D}$ for 1000 random realizations of $\hat{H}_{\mathrm{nh}}$ by choosing its eigenvalues $\{\lambda_{m}\}$ from a Gaussian distribution with zero mean and unit variance. In Fig. 1(a), we plot the maximum as well as minimum among all the random realizations, denoted by $\mathcal{D}_{\max}$ and $\mathcal{D}_{\min}$, respectively, against $\bar{\alpha}=1-\alpha$. At $\alpha=1$ ($\bar{\alpha}=0$), the right and left eigenvectors recover perfect Bell states and the Hamiltonian becomes Hermitian, reflected in vanishing $\mathcal{D}$. Though fluctuating, $\mathcal{D}_{\min}$ tends to increase with $\bar{\alpha}$. The growth of $\mathcal{D}_{\max}$ is smoother and at a faster rate, reaching the maximum non-Hermiticity of $\mathcal{D}=1$ at $\bar{\alpha}=1$.
@@ III.1.2 Non-Hermiticity score for single-site property: magnetization
One of the main goals of our work is to examine to what extent the non-Hermiticity of the model Hamiltonian gets translated to its physical properties. Towards that aim, we now analyze the non-Hermiticity score for the magnetization in individual eigenstates. Results for quantum entanglement in individual eigenstates are presented in the following section. Analytical forms of all the quantities discussed in these two sections are provided in Appendix D. Two subsequent sections contain results for purity and entanglement under time evolution.
In Fig. 2(a), we plot the non-Hermiticity score for the magnetization of the first qubit (labelled $A$) along the $z$-axis, $\mathcal{F}[\rho]=m_{z}[\rho]=\text{Tr}(\hat{\sigma}^{z}_{A}\rho)$, $SC^{m_{z}}_{k}=|m_{z}[\rho_{RR}^{k}]-m_{z}[\rho_{LL}^{k}]|$ Here, $\hat{\sigma}^{z}$ is the $2\times 2$ Pauli matrix along the $z$-direction and $\rho_{RR}^{k}=\frac{|R_{k}\rangle\langle R_{k}|}{\langle R_{k}|R_{k}\rangle}$, $\rho_{LL}^{k}=\frac{|L_{k}\rangle\langle L_{k}|}{\langle L_{k}|L_{k}\rangle}$, where $|R_{k}\rangle$ and $|L_{k}\rangle$ are individual right and left eigenvectors of $H^{\mathrm{Bell}}_{\mathrm{nh}}$, as given in Eqs. (4) and (5). As we are considering here properties of single eigenvectors, the eigenvalues $\lambda_{k}$ do not directly enter the non-Hermiticity score, in contrast to $\mathcal{D}$ discussed above, and no sampling over $\lambda_{k}$ is necessary. To give an operational interpretation of the above observables, we may want to assume the eigenvalues to fulfill case 2 in Sec. C, such that $\rho_{RR}^{k}$ and $\rho_{LL}^{k}$ are prepared when evolving an initial state to long times under $\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}}$ and $(\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}})^{\dagger}$, respectively.
@@ III.1.3 Non-Hermiticity score for the von Neumann entropy (static case)
In this section, we analyze entanglement properties of the same set of right and left eigenvectors as considered above, in particular of the von Neumann entropy (VNE), defined as $)\mathcal{S}[\hat{\rho}]=-\text{Tr}_{A}(\text{Tr}_{B}(\hat{\rho})\log_{2}\text{Tr}_{B}(\hat{\rho}$)). Figure 2(b) presents the associated non-Hermiticity score, $SC^{\mathcal{\mathcal{S}}}_{k}=|\mathcal{S}[\hat{\rho}_{RR}^{k}]-\mathcal{S}[\hat{\rho}_{LL}^{k}]|$. Unlike the case for $m_{z}$, at each value of $\bar{\alpha}$, $SC^{\mathcal{S}}_{k}$ becomes minimum for $k=1,2$, and the behavior does not comply with $\mathcal{D}$. In particular, $SC^{\mathcal{S}}_{k=1,2}$ vanishes both for $\bar{\alpha}\rightarrow 0$ and $\bar{\alpha}\rightarrow 1$, and attains a maximum around $\bar{\alpha}=0.449$. This is significantly different from the behavior of $\mathcal{D}_{\max}$ (or $\mathcal{D}_{\min}$) as well as the single-site property discussed above.
In contrast, the maximum value of $SC^{\mathcal{S}}_{k}$, attained for all values of $\bar{\alpha}$ at $k=3,4$, remains akin to $\mathcal{D}$. Nevertheless, quantitative differences remain. E.g., unlike $\mathcal{D}$, $SC^{\mathcal{S}}_{k=3,4}$ increases almost linearly in the region $0.4\lesssim\bar{\alpha}\leq 1.0$ and goes to zero at $\bar{\alpha}=0$. Moreover, $SC^{\mathcal{S}}_{k=3,4}$ does not reach its theoretical maximum of $1$ at $\bar{\alpha}=1$. Again, the VNE obtained for the biorthogonal ensemble $\hat{\rho}_{RL}^{k}=|R_{k}\rangle\langle L_{k}|$ remains independent of $\bar{\alpha}$ (yielding the value $\mathcal{S}(|R_{k}\rangle\langle L_{k}|)=\frac{1}{2}$ $\forall k$) and is thus not a useful probe for the non-Hermiticity of the system.
@@ III.1.4 Non-Hermiticity score of global purity
In this section, we study the time-dependent non-Hermiticity score of the purity, $\mathcal{P}[\hat{\rho}(t)]={\text{Tr}\Big(\hat{\rho}(t)^{2}\Big)}/{\text{Tr}(\hat{\rho}(t))^{2}}$, of a state $\hat{\rho}(t)$ obtained from an initial state $\hat{\rho}_{\mathrm{in}}$ when evolved under $\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}}$ or $(\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}})^{\dagger}$. If the initial state is chosen as maximally mixed, $\hat{\rho}_{\mathrm{in}}=\frac{\mathbb{I}}{4}$, one gets $\mathcal{P}(\hat{\rho}_{RR}(t))=\mathcal{P}(\hat{\rho}_{LL}(t))$ (see Appendix E). Hence, $SC^{\mathcal{P}}(t)=0$ at all times and fails to witness any imprint of $\mathcal{D}$. However, for a non-maximally mixed initial state, $SC^{\mathcal{P}}(t)$ acquires a non-zero value. In Fig. 3, we plot the time-dependent purity score for systems initialized in a mixed Werner state, $\hat{\rho}_{W}=\delta|\Psi^{-}\rangle\langle\Psi^{-}|+(1-\delta)\frac{\mathbb{I}}{4}$, with $|\Psi^{-}\rangle=\frac{|01\rangle-|10\rangle}{\sqrt{2}}$. We tune the initial purity to (a) $\delta=0.1$, (b) $\delta=0.5$, and (c) $\delta=0.9$, and choose four different values of $\alpha$, $\alpha=0.1$, $0.3$, $0.6$, $0.9$, which are depicted using lighter to darker shades of blue. In these plots, we fix the Hamiltonian eigenvalues to $\lambda_{1}=0.1$, $\lambda_{2}=0.2$, $\lambda_{3}=0.3$, and $\lambda_{4}=0.4$.
@@
Figure 3: Behavior of non-Hermiticity score with time, for global purity $\mathcal{P}$ and von Neumann entropy $\mathcal{S}$, computed for an initial state $\hat{\rho}_{W}=\delta|\Psi^{-}\rangle\langle\Psi^{-}|+(1-\delta)\frac{\mathbb{I}}{4}$ subjected to evolution under $\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}}$ and $(\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}})^{\dagger}$. We probe the cases for $\alpha$=0.1, 0.3, 0.6, 0.9 as shown using curves with darker to lighter shades of blue, and in all cases fix the Hamiltonian eigenvalues to $\lambda_{1}=0.1,\ \lambda_{2}=0.2,\ \lambda_{3}=0.3,\ \lambda_{4}=0.4$. (a)-(c) Purity score, $SC^{\mathcal{P}}(t)$ for (a) $\delta=0.1$, (b) $\delta=0.5$, and (c) $\delta=0.9$. For large $\delta$, the amplitude of $SC^{\mathcal{P}}(t)$ decreases with increasing $\alpha$, while for small $\delta$ it attains a maximum at intermediate values of $\alpha$. (d)-(f) Non-Hermiticity score for the VNE, $SC^{\mathcal{S}}(t)$, for same choices of $\delta$.